OpenStudy (anonymous):

7x + 10y = 37 4x + 5y = 19

OpenStudy (anonymous):

you may solve by elimination method: multiply equation 2 by -2; make linear combination with equation 1 7x + 10y = 37 -8x + -10y = -38 ============ -x =-1 x=1 now sub x=1 into any equation above to find y; I choose the original equation 2 4(1) + 5y = 19 (sub x=1) 5y = 19-4 (subtract 4) y=3 (divide by 5) the solution is x=1; y=3, or you may write it as an ordered pair (1, 3)